Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 120,000 liters at a budgeted price of $225 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials Direct labor (2 pounds @ $14) (0.5 hours @ $44) $28 22 Variable overhead is applied based on direct labor hoPlease show all of the steps, thank you very much.
Answer #1

a. Direct Material

Price Variance = (standard price - actual price ) X actual quantity used.

($14 - $14.2244) X 196000 pounds

= $44000 ( unfavourable , since answer is in negative )

work note:

Actual price of material

$2788000 total materials cost / 196000 pounds of materials purchased }

Material Efficiency variance = (Actual unit usage - Standard unit usage) x Standard cost per unit

Working note - Standard unit usage for actual output

Actual output - work note.

Actual sales value = $ 25938000

assumption. : standard sales price per unit  and actual sale price per unit is same.

Actual production will be = 25938000/225 = 115280 litres

Standard materials requirement for actual production now will be.

115280 litres X 2 pounds = 230560 pounds

Now applying formula.

(Actual unit usage - Standard unit usage) x Standard cost per unit

(196000 - 230560 ) X 14 = $483,840 unfavourable

b. Labour variances:

labour price variance

(Actual Hours x Actual Rate   - Actual Hours x Standard Rate)

Actual hours = 50400 hours X actual rate 49.8015 = $2510000

work note : actual rate = direct labour cost / direct labour hours

= 2510000/ 50400 = 49.8015

Actual Hours x Standard Rate = 50400 X44 =$2217600

labour price variance = $292,400 unfavourable

Direct labour efficiency variance = (Actual Hours x Standard Rate -Standard Hours x Standard Rate)

50400 X 44 - standard hours for actual production X 44

work note : standard hours for actual production

= actual production 115280 X 0.5 hours = 57640 hours

therefore = 50400 X 44 - 57640 X 44

=2217600- 2536160

=$ 318,560 favourable

c. variable overhead

price variance = (Actual Hours x Actual Rate   - Actual Hours x Standard Rate)

($6480000 - 50400 x 120) = $432000 unfavourable

efficiency variance.

(Actual hours - Standard hours) x Standard rate = variable overhead efficiency variance

(50400- 57640)X120 = $868,800 favourable

Direct Material
Price Variance $44000 U
Material Efficiency variance $483,840 U
Direct Labour
Price Variance $292,400 U
Efficiency Variance $ 318,560 F
Variable overhead
Price Variance $432,000 U
Efficiency Variance $868,800 F